# Nonlinearly charging the conformally dressed black holes preserving duality and conformal invariance

### v1

@@ IV Nonlinearly charging the MTZ black hole via ModMax
We now proceed to charge the conformally dressed MTZ black holes [24, 25] beyond the linear regime but still preserving the conformal and duality symmetries of its involved sources. In concrete, this means considering the ModMax nonlinear electrodynamics [30] instead of Maxwell, and hence dealing with the following action principle
@@
We are interested in constructing at once stationary solutions extended with a Taub-NUT parameter. This goal is motivated by the fact that on the one hand, the NUT solution generalizing the conventionally charged conformally dressed black hole is already known [27], and on the other, we can exploit the fact that the NUT solutions were generically obtained for all duality-invariant nonlinear electrodynamics in absence of the scalar field [31]. Regarding the latter, the precise case of ModMax was recently re-analyzed in Refs. [40, 41]. It is easy to incorporate a constant conformal scalar field on top of the previous configurations [42], but the real challenge is to achieve the goal with a nontrivial scalar field. All the previously mentioned configurations are obviously inspired by the vacuum spacetime solutions originally found and characterized in Refs. [43, 44, 45]. Starting from this observation, we will consider the following generic Taub-NUT ansatz for our task
@@
determines the vector potentials with different parameterizations. Here, $m$, $q$, and $p$ are integration constants that together with $n$ are tied via the coupling constants of the theory through the following relation
@@
A careful examination of the solution reveals that, in units where $\kappa=8\pi$, $m$ is exactly the mass of the spherical black holes resulting at $k=1$ in the static limit $n=0$ we discuss later; this can be calculated using the successful off-shell formalism proposed in [46] and discussed in detail in [47], see also [48]. Besides, $q$ and $p$ are nothing but the electric and magnetic charges, respectively, which is explicitly checked also in the spherical case $k=1$ by integrating the right-hand sides of (6) as usual over a sphere at infinity.
@@
giving for the Euclidean vector potentials $A_{t}(r)=h(r;p,q)$ and $A^{*}_{t}(r)=h(r;q,p)$. The (anti-)self-dual points are defined by $m=\pm n$ and $p=\pm q$, yielding $A=\pm A^{*}$ with vanishing electromagnetic energy-momentum tensor and making the scalar field trivial. Thus reducing the Euclidean metric to a vacuum configuration with (anti-)self-dual curvature, in agreement with previous results [40, 41, 49, 50].
Before continuing with the characterization of the solutions, we want to ponder on the meaning of the constraint (38f). Indeed, it has long been known that the Weyl tensor decomposition in terms of electric and magnetic parts in the paradigm of Taub-NUT spacetimes provides gravito-electric and gravito-magnetic interpretations for the mass $m$ and the NUT parameter $n$, respectively [51]. In this sense the dyonic character of solutions like (38) is twofold, since it is not only described by the two electromagnetic charges but also by the two gravito-electromagnetic peers. This provides a remarkable interpretation of the twofold dyonic version of restriction (38f) as a proportionality relation, mostly via the coupling constants, between a sort of total dyonic electromagnetic charge $q^{2}+p^{2}$ and its dyonic gravitational analog $m^{2}+n^{2}$, i.e. establishing a correlation between both dyonic counterparts.
@@
generalizing to the dyonic case the purely electric, $p=0$, conformal Taub-NUT solution of Ref. [27]. The corresponding metric and scalar field are unaffected by those limits, which brings up the following interesting curiosity of solution (38). On the one hand, the metric and scalar field behave exactly the same way as in the linearly charged case of [27]. On the other hand, the electromagnetic fields of (38) are exactly the same as those of ModMax without a scalar field [40, 41]. This peculiarity is a direct consequence of the decoupling between both kinds of involved conformal matter and additionally, has roots in that the constitutive relations (36a) turn out to have no information on the particular gravitational potential of the background (37a). In fact, only the constraint (38f) non-trivially merges all the involved ingredients. This is in analogy with what occurs in the linear case when the neutral topological MTZ black holes are conventionally charged [24, 25].
@@
and its purely electric (or magnetic) limit would imply $\mathscr{Q}=0$. Thus, in order to ensure that the new black holes (44) probes the fully nonlinear regime it is mandatory to consider dyonic configurations, as those obtained in Ref. [52] in the absence of the scalar field.
@@ V New (super-)renormalizably dressed spacetimes
The aim of this section is to generalize the conformally dressed spacetimes (38) to new (super-)renormalizably dressed examples by exploiting the results already established in Ref. [28]. In order to accomplish this goal, we need to extend first the above map to other conformally charged configurations not necessarily rigged by the linearity of Maxwell theory, and in particular to those charged by the ModMax electrodynamics where in general $\gamma\neq 0$ in action (35a). Since everything works by scaling the Maxwell action as $1-a^{2}$, it is enough to demand that under the sought transformation the same occurs for any conformal electromagnetic action. In doing so, we recall that the structural functions of every conformally invariant nonlinear electrodynamics are necessarily homogeneous, as was explained in Subsec. III.2. Concretely, let us remember that the degree of the “Hamiltonian” as a function of the standard invariants must be one, i.e. it satisfies $\mathscr{H}(\hat{\lambda}\mathscr{P},\hat{\lambda}\mathscr{Q})=\hat{\lambda}\mathscr{H}(\mathscr{P},\mathscr{Q})$. Hence, it is clear that the appropriate transformation for the second vector potential, generalizing the map of Ref. [28] to include any conformal electrodynamics, must be necessarily given by the same scaling law
@@
where $\bar{R}$ refers to the scalar curvature associated with the metric $\bar{g}$. Additionally, we have defined $\bar{F}=\textrm{d}\bar{A}$ and ¯$\bar{\star}\bar{P}=\textrm{d}\bar{A}^{*}$, accordingly $\bar{L}_{\text{MM}}$ is the ModMax Lagrangian (35c) for these fields. At the same time, the conformal scalar Lagrangian changes to
@@
with $\partial\Sigma$ consequently taken as a sphere at infinity. It is worth mentioning that for purely electric configurations, $p=0$, of the Maxwell case, $\gamma=0$, the above solution becomes that of Ref. [29]. Consequently, setting $a=0$ in the (super-)renormalizably dressed solution (50) one recovers the conformally dressed seed (38). The other way around, taking as seeds all the special limits of (38) —namely (43), (44), and (46)— new special (super-)renormalizably dressed solutions are straightforwardly obtained, since those limits commute with the map there is no need to explicitly write them.
@@ VI Nonlinearly charging the non-Noetherian conformal sector
It is certainly known now that the most general second-order equation describing a conformally invariant scalar field that is extreme of an action principle is definitely not Eq. (36b), as it has been recently proved in [19] inspired in the previous work of [20]. In particular, it was originally identified in Ref. [20] that a non-Noetherian conformal contribution can be also considered in Eq. (36b). It comes from the following additional piece to Lagrangian (35b) that extends the action principle (35a) used in Sec. IV to
@@
where $\mathscr{G}=R^{2}-4R_{\alpha\beta}R^{\alpha\beta}+R_{\alpha\beta\mu\nu}R^{\alpha\beta\mu\nu}$ is the Gauss-Bonnet density, $G^{\mu\nu}$ the Einstein tensor, and $\Box$ the d’Alembert operator, being all defined with the standard metric $g_{\mu\nu}$. With this new contribution, Eq. (36b) is extended as
@@
where $\tilde{\mathscr{G}}$ is the Gauss-Bonnet density again, but now that corresponding to the auxiliary metric $\tilde{g}_{\mu\nu}=\Phi^{2}g_{\mu\nu}$, which by construction is invariant under conformal transformations
@@
Since the new term with tilde in (53) behaves as a constant under conformal transformations, the resulting equation is again conformally invariant. However, the action built from Lagrangian (52b) is not preserved under the transformations (54) and consequently, the conformal invariance of (53) is not due to the Noether theorem. This is why Ref. [19] termed non-Noetherian conformal scalar fields to those whose dynamic is governed by the previous equation.
As an Euler-Lagrange equation of a related action, the second-order conformally invariant properties of (53) are not spoiled by adding to Lagrangian (52b) arbitrary nonminimal couplings to gravity build with the Weyl tensor of the auxiliary metric $\tilde{g}_{\mu\nu}$ [19]. Nevertheless, only the extension (52b) gives rise to an also second-order addition to the energy-momentum tensor (36d), since the Lagrangian (52b) remarkably belongs to the most general family of second-order variational principles for the scalar field and the metric which are defined by the Horndeski action [53]. We do not write here the involved explicit expression of the new contribution to the energy-momentum tensor for the economy of space, but it can be consulted in Ref. [20].
The static spherically-symmetric non-Noetherian conformally dressed black holes, charged with the $\gamma=0$ Maxwell sector of action (52), were originally derived in Ref. [20], where Fernandes shows the existence of two independent branches of solutions. There is no known Taub-NUT generalization of the Fernandes static branches. Hence, here we focus on the solutions of the non-Noetherianly extended conformal system derived from action (52) in the static limit, $n=0$, of our study ansatz (37). When the ModMax electrodynamics is used to nonlinearly charge the two Fernandes branches of non-Noetherian conformal black holes beyond the Maxwell regime, such branches are generalized as follows
@@
For $\gamma=0$ and $k=1$ the conventionally-charged spherically-symmetric Fernandes black holes [20], dressed with a non-Noetherian conformal scalar field, are recovered. In order to ease the comparison we make transparent the relation between the Fernandes notations and ours
@@
We found it appropriate to emphasize the further bifurcation (57b) of the second branch not reported in Ref. [20], since the spherical $\alpha<0$ sector originally found in such reference is not extensible to flat $k=0$ or hyperbolic $k=-1$ topologies, which are only covered by the new $\alpha>0$ sector. In fact, the latter is the only non-Noetherian conformally dressed solution at all surviving in the limit $k\to 0$. This is compatible with the recent results of Refs. [54] and [55] for more general, non-necessarily symmetric, Horndeski theories allowing a similar branching structure to the one first found by Fernandes in Ref. [20]. There, generic planar solutions $k=0$ unavoidably require the absence of the standard Noetherian conformal sector $L_{\text{CS}}$ (35b), which is impossible here. Besides, if the limit $\alpha=0$ is assumed then the metric lower-sign solution becomes the ModMax nonlinearly charged black holes of [52] generalizing the dyonic Reissner-Nordström ones, where the scalar field is absent.
@@ VII Concluding Remarks
In this work, we have considered self-gravitating configurations of ModMax nonlinear electrodynamics together with a self-interacting scalar field conformally coupled to gravity. The distinctive feature of ModMax theory is that it is the unique nonlinear extension of the Maxwell theory enjoying its conformal invariance as well as its duality symmetry [30]. We have taken advantage of this work to present a novel and transparent derivation of this electrodynamics by employing the dual Legendre formulations introduced in Ref. [31]. Of particular relevance is their simple characterization of duality invariant theories in contrast with that allowed for the standard Lagrangian formalism [37]. Our approach demands that the Legendre duals themselves be invariant under duality and conformal transformations. In doing so, we end up with elementary linear first-order equations which straightforwardly yield as unique solution the ModMax electrodynamics [30].
Correspondingly, a limitation of our work is the same as the approach that we borrow from Ref. [31] to duality invariance, i.e. it only applies to the most common algebraically general electromagnetic configurations where not all the eigenvalues vanish [32], making sense of the alignment (8). This means that algebraically special or null electromagnetic fields, where all the eigenvalues are identically zero [32], are not included in our analysis. Therefore, null fields can still give rise to potentially different realizations of duality invariance excluded from those treated here. This seems to be the case of those configurations rigged by the Białynicki-Birula electrodynamics that is invariant under $SL(2,\mathbb{R})$ transformations more than duality rotations and additionally also respects conformal symmetry [56, 57].
Regarding the ModMax nonlinearly charged spacetimes, our most general result is the family of solutions (50), which describes non-static configurations whose background is dyonically charged by ModMax fields and is simultaneously dressed by a conformally coupled and (super-)renormalizable scalar field. Algebraically, the spacetime is classified in the Petrov scheme as of type $D$. Whereas, geometrically, it is conformally related to a complex line bundle fibered on a Kähler manifold. As such, the solution is closely related to the classical Taub-NUT vacuum metric [43, 44, 45]. In the static limit, the fibration is trivialized and also black hole solutions emerge. In that case, the resulting black holes are nonlinearly charged generalizations of those presented in Ref. [28], which in turn has as neutral limits the ones found in Ref. [26].
@@
We expect the emergence of all the presented solutions definitely encourages further research into the fertile subject of scalar fields conformally coupled to gravity and supplemented by nonlinear electrodynamics, especially in light of the recent developments discussed in Refs. [49, 58, 59, 50]. A very promising path within all the potential possibilities is the recent better understanding of accelerating NUT spacetimes in General Relativity. Thanks to new parameterizations [60, 61] of the celebrated Plebański-Demiański solutions [62], it is now certainly understood that the absence of rotation excludes the simultaneous presence of acceleration and NUT parameter in Petrov type D spacetimes. This finally clarifies the puzzle of why the accelerated NUT vacuum spacetimes originally found in [63] are outside this special class, being in fact of general Petrov type [64]. The connection with our work is that on the one hand, the accelerated NUT spacetimes have been recently extended to be sourced by free conformal scalar fields [65], and on the other hand, the first nonlinearly charged accelerated spacetimes were precisely built with ModMax [66].

### v2

@@ IV Nonlinearly charging the MTZ black hole via ModMax
We now proceed to charge the conformally dressed MTZ black holes [25, 26] beyond the linear regime but still preserving the conformal and duality symmetries of its involved sources. In concrete terms, this means considering the ModMax nonlinear electrodynamics [31] instead of Maxwell, and hence dealing with the following action principle
@@
We are interested in constructing at once stationary solutions extended with a Taub-NUT parameter. This goal is motivated by the fact that, on the one hand, the NUT solution generalizing the conventionally charged conformally dressed black hole is already known [28], and on the other, we can exploit the fact that the NUT solutions were generically obtained for all duality-invariant nonlinear electrodynamics in the absence of the scalar field [32]. Regarding the latter, the precise case of ModMax was recently reanalyzed in Refs. [40, 41]. It is easy to incorporate a constant conformal scalar field on top of the previous configurations [42], but the real challenge is to achieve the goal with a nontrivial scalar field. All the previously mentioned configurations are obviously inspired by the vacuum spacetime solutions originally found and characterized in Refs. [43, 44, 45]. Starting from this observation, we will consider the following generic Taub-NUT ansatz for our task
@@
determines the vector potentials with different parameterizations. Here, $m$, $q$, and $p$ are integration constants that, together with $n$, are tied via the coupling constants of the theory through the following relation
@@
A careful examination of the solution reveals that, in units where $\kappa=8\pi$, $m$ is exactly the mass of the spherical black holes resulting at $k=1$ in the static limit $n=0$ we discuss later. This can be calculated using the successful off-shell formalism proposed in [46] and discussed in detail in [47]; see also [48]. Besides, $q$ and $p$ are nothing but the electric and magnetic charges, respectively, which is explicitly checked also in the spherical case $k=1$ by integrating the right-hand sides of (6) as usual over a sphere at infinity.
@@
giving for the Euclidean vector potentials $A_{t}(r)=h(r;p,q)$ and $A^{*}_{t}(r)=h(r;q,p)$. The (anti-)self-dual points are defined by $m=\pm n$ and $p=\pm q$, yielding $A=\pm A^{*}$ with vanishing electromagnetic energy-momentum tensor and making the scalar field trivial, thus reducing the Euclidean metric to a vacuum configuration with (anti-)self-dual curvature, in agreement with previous results [40, 41, 49, 50].
Before continuing with the characterization of the solutions, we want to ponder the meaning of the constraint (38f). Indeed, it has long been known that the Weyl tensor decomposition in terms of electric and magnetic parts in the paradigm of Taub-NUT spacetimes provides gravitoelectric and gravitomagnetic interpretations for the mass $m$ and the NUT parameter $n$, respectively [51]. In this sense the dyonic character of solutions like (38) is twofold, since it is not only described by the two electromagnetic charges but also by the two gravitoelectromagnetic peers. This provides a remarkable interpretation of the twofold dyonic version of restriction (38f) as a proportionality relation, mostly via the coupling constants, between a sort of total dyonic electromagnetic charge $q^{2}+p^{2}$ and its dyonic gravitational analog $m^{2}+n^{2}$, i.e. establishing a correlation between both dyonic counterparts.
@@
generalizing to the dyonic case the purely electric, $p=0$, conformal Taub-NUT solution of Ref. [28]. The corresponding metric and scalar field are unaffected by those limits, which brings up the following interesting curiosity of solution (38). On the one hand, the metric and scalar field behave exactly the same way as in the linearly charged case of [28]. On the other hand, the electromagnetic fields of (38) are exactly the same as those of ModMax without a scalar field [40, 41]. This peculiarity is a direct consequence of the decoupling between both kinds of involved conformal matter and additionally, has roots in that the constitutive relations (36a) turn out to have no information on the particular gravitational potential of the background (37a). In fact, only the constraint (38f) nontrivially merges all the involved ingredients. This is in analogy with what occurs in the linear case when the neutral topological MTZ black holes are conventionally charged [25, 26].
@@
and its purely electric (or magnetic) limit would imply $\mathscr{Q}=0$. Thus, in order to ensure that the new black holes (44) probe the fully nonlinear regime it is mandatory to consider dyonic configurations, as those obtained in Ref. [52] in the absence of the scalar field.
@@ V New (super-)renormalizably dressed spacetimes
The aim of this section is to generalize the conformally dressed spacetimes (38) to new (super-)renormalizably dressed examples by exploiting the results already established in Ref. [29]. In order to accomplish this goal, we need to extend first the above map to other conformally charged configurations not necessarily rigged by the linearity of Maxwell theory. In particular, to those charged by the ModMax electrodynamics where $\gamma\neq 0$ in action (35a). Since everything works by scaling the Maxwell action as $1-a^{2}$, it is enough to demand that under the sought transformation the same occurs for any conformal electromagnetic action. In doing so, we recall that the structural functions of every conformally invariant nonlinear electrodynamics are necessarily homogeneous, as was explained in Sec. III.2. Concretely, let us remember that the degree of the Hamiltonian as a function of the standard invariants must be one; i.e., it satisfies $\mathscr{H}(\hat{\lambda}\mathscr{P},\hat{\lambda}\mathscr{Q})=\hat{\lambda}\mathscr{H}(\mathscr{P},\mathscr{Q})$. Hence, it is clear that the appropriate transformation for the second vector potential, generalizing the map of Ref. [29] to include any conformal electrodynamics, must be necessarily given by the same scaling law
@@
where $\bar{R}$ refers to the scalar curvature associated with the metric $\bar{g}$. Additionally, we have defined $\bar{F}=\textrm{d}\bar{A}$ and ¯$\bar{\star}\bar{P}=\textrm{d}\bar{A}^{*}$; accordingly, $\bar{L}_{\text{MM}}$ is the ModMax Lagrangian (35c) for these fields. At the same time, the conformal scalar Lagrangian changes to
@@
with $\partial\Sigma$ consequently taken as a sphere at infinity. It is worth mentioning that for purely electric configurations, $p=0$, of the Maxwell case, $\gamma=0$, the above solution becomes that of Ref. [30]. Consequently, setting $a=0$ in the (super-)renormalizably dressed solution (50) one recovers the conformally dressed seed (38). The other way around, taking as seeds all the special limits of (38)—namely (43), (44), and (46)—new special (super-)renormalizably dressed solutions are straightforwardly obtained; since those limits commute with the map there is no need to explicitly write them.
@@ VI Nonlinearly charging the non-Noetherian conformal sector
It is certainly known now that the most general second-order equation describing a conformally invariant scalar field that is the extreme of an action principle is definitely not Eq. (36b), as it has been recently proved in [20] inspired in the previous work of [21]. In particular, it was originally identified in Ref. [21] that a non-Noetherian conformal contribution can be also considered in Eq. (36b). It comes from the following additional piece to Lagrangian (35b) that extends the action principle (35a) used in Sec. IV to
@@
where $\mathscr{G}=R^{2}-4R_{\alpha\beta}R^{\alpha\beta}+R_{\alpha\beta\mu\nu}R^{\alpha\beta\mu\nu}$ is the Gauss-Bonnet density, $G^{\mu\nu}$ the Einstein tensor, and $\Box$ the d’Alembert operator, all being defined with the standard metric $g_{\mu\nu}$. With this new contribution, Eq. (36b) is extended as
@@
where $\tilde{\mathscr{G}}$ is the Gauss-Bonnet density again, but now corresponding to the auxiliary metric $\tilde{g}_{\mu\nu}=\Phi^{2}g_{\mu\nu}$, which by construction is invariant under conformal transformations
@@
Since the new term with the tilde in (53) behaves as a constant under conformal transformations, the resulting equation is again conformally invariant. However, the action built from Lagrangian (52b) is not preserved under the transformations (54) and consequently, the conformal invariance of (53) is not due to the Noether theorem. This is why Ref. [20] termed non-Noetherian conformal scalar fields to those whose dynamic is governed by the previous equation. [^2] For the two-dimensional non-Noetherian analog see Refs. [53, 54].
As an Euler-Lagrange equation of a related action, the second-order conformally invariant properties of (53) are not spoiled by adding to Lagrangian (52b) arbitrary nonminimal couplings to gravity build with the Weyl tensor of the auxiliary metric $\tilde{g}_{\mu\nu}$ [20]. Nevertheless, only the extension (52b) also gives rise to a second-order addition to the energy-momentum tensor (36d), since the Lagrangian (52b) remarkably belongs to the most general family of second-order variational principles for the scalar field and the metric which are defined by the Horndeski action [55]. We do not write here the involved explicit expression of the new contribution to the energy-momentum tensor for the economy of space, but it can be consulted in Ref. [21].
The static spherically symmetric non-Noetherian conformally dressed black holes, charged with the $\gamma=0$ Maxwell sector of action (52), were originally derived in Ref. [21], where Fernandes showed the existence of two independent branches of solutions. There is no known Taub-NUT generalization of the Fernandes static branches. Hence, here we focus on the solutions of the non-Noetherianly extended conformal system derived from action (52) in the static limit, $n=0$, of our study ansatz (37). When the ModMax electrodynamics is used to nonlinearly charge the two Fernandes branches of non-Noetherian conformal black holes beyond the Maxwell regime, such branches are generalized as follows
@@
For $\gamma=0$ and $k=1$ the conventionally charged spherically symmetric Fernandes black holes [21], dressed with a non-Noetherian conformal scalar field, are recovered. In order to ease the comparison we make transparent the relation between the Fernandes notations and ours
@@
We found it appropriate to emphasize the further bifurcation (57b) of the second branch not reported in Ref. [21], since the spherical $\alpha<0$ sector originally found in that reference is not extensible to flat $k=0$ or hyperbolic $k=-1$ topologies, which are only covered by the new $\alpha>0$ sector. In fact, the latter is the only non-Noetherian conformally dressed solution at all surviving in the limit $k\to 0$. This is compatible with the recent results of Refs. [56] and [57] for more general, non-necessarily symmetric, Horndeski theories allowing a similar branching structure to the one first found by Fernandes in Ref. [21]. There, generic planar solutions $k=0$ unavoidably require the absence of the standard Noetherian conformal sector $L_{\text{CS}}$ (35b), which is impossible here. Besides, if the limit $\alpha=0$ is assumed then the metric lower-sign solution becomes the ModMax nonlinearly charged black holes of [52] generalizing the dyonic Reissner-Nordström ones, where the scalar field is absent.
@@ VII Concluding Remarks
In this work, we have considered self-gravitating configurations of ModMax nonlinear electrodynamics together with a self-interacting scalar field conformally coupled to gravity. The distinctive feature of ModMax theory is that it is the unique nonlinear extension of the Maxwell theory enjoying its conformal invariance as well as its duality symmetry [31]. We have taken advantage of this work to present a novel and transparent derivation of this electrodynamics by employing the dual Legendre formulations introduced in Ref. [32]. Of particular relevance is their simple characterization of duality-invariant theories in contrast with that allowed for the standard Lagrangian formalism [38]. Our approach demands that the Legendre duals themselves be invariant under duality and conformal transformations. In doing so, we end up with elementary linear first-order equations which straightforwardly yield as unique solution the ModMax electrodynamics [31].
Correspondingly, a limitation of our work is the same as the approach that we borrow from Ref. [32] to duality invariance, i.e. it only applies to the most common algebraically general electromagnetic configurations where not all the eigenvalues vanish [33], making sense of the alignment (8). This means that algebraically special or null electromagnetic fields, where all the eigenvalues are identically zero [33], are not included in our analysis. Therefore, the use of different variables including the particular case of null fields can still give rise to potentially different realizations of duality invariance excluded from those treated here. This seems to be the case of those configurations rigged by the Białynicki-Birula electrodynamics that is invariant under $SL(2,\mathbb{R})$ transformations more than duality rotations and additionally also respects conformal symmetry [58, 59].
Regarding the ModMax nonlinearly charged spacetimes, our most general result is the family of solutions (50), which describes nonstatic configurations whose background is dyonically charged by ModMax fields and is simultaneously dressed by a conformally coupled and (super-)renormalizable scalar field. Algebraically, the spacetime is classified in the Petrov scheme as of type $D$. Whereas, geometrically, it is conformally related to a complex line bundle fibered on a Kähler manifold. As such, the solution is closely related to the classical Taub-NUT vacuum metric [43, 44, 45]. In the static limit, the fibration is trivialized and also black hole solutions emerge. In that case, the resulting black holes are nonlinearly charged generalizations of those presented in Ref. [29], which in turn has as neutral limits the ones found in Ref. [27].
@@
We expect the emergence of all the presented solutions definitely encourages further research into the fertile subject of scalar fields conformally coupled to gravity and supplemented by nonlinear electrodynamics, especially in light of the recent developments discussed in Refs. [49, 60, 61, 50]. A very promising path within all the potential possibilities are new results [62, 63, 64, 65, 66] on the celebrated Plebański-Demiański solutions [67], since thanks to them the presence of acceleration and NUT parameter in Petrov type D spacetimes is now better understood. This finally clarifies the puzzle of why the accelerated NUT vacuum spacetimes originally found in [68] are outside this special class, being in fact of general Petrov type [69]. The connection with our work is that on the one hand, the accelerated NUT spacetimes have been recently extended to be sourced by free conformal scalar fields [70], and on the other hand, the first nonlinearly charged accelerated spacetimes were precisely built with ModMax [71].